# Continuous sentiment scores for literary and multilingual contexts

### v1

@@ Abstract
Sentiment Analysis is widely used to quantify sentiment in text, but its application to literary texts poses unique challenges due to figurative language, stylistic ambiguity, as well as sentiment evocation strategies. Traditional dictionary-based tools often underperform, especially for low-resource languages, and transformer models, while promising, typically output coarse categorical labels that limit fine-grained analysis. We introduce a novel continuous sentiment scoring method based on concept vector projection, trained on multilingual literary data, which more effectively captures nuanced sentiment expressions across genres, languages, and historical periods. Our approach outperforms existing tools on English and Danish texts, producing sentiment scores whose distribution closely matches human ratings, enabling more accurate analysis and sentiment arc modeling in literature.
@@ 1 Introduction & Related Works
Sentiment analysis quantifies sentiment in text and is widely used across domains, from product reviews analysis to social media monitoring [32, 7]. In computational literary studies, it is increasingly employed to model narrative dynamics, particularly for modelling sentiment arcs across novels [16, 26, 38, 6]. This typically requires continuous sentiment scores, mapping sentiment along a spectrum rather than using categorical labels, such as positive/negative.
Despite the growing use of continuous sentiment scoring in literary studies, the validity of current tools in capturing literary sentiment expression remains underexplored. Popular tools such as Syuzhet have faced severe criticism for oversimplification or poor generalizability [31] – issues that point to broader limitations in applying off-the-shelf sentiment analysis tools to literary texts. Indeed, the literary domain poses distinct challenges: figurative language, multiple narrative layers, and stylistic ambiguity all complicate sentiment detection [8, 3].
More recent transformer-based models appear better equipped to handle the complexity of literary language [30], and techniques that transform categorical model outputs into continuous scores have been proposed [4]. Such methods have proven more effective than tailored dictionary-based tools, particularly in low-resource language settings and across languages [13]. However, empirical benchmarks comparing model predictions to human judgments remain limited, particularly in languages other than English.
@@
    Cross-lingual performance drops. Most Sentiment Analysis tools are developed for high-resource languages, and their transfer to low-resource ones like Danish is non-trivial. Although Danish has several dictionary-based tools (i.a., Asent [12], Sentida [21]), these have seen little use on historical literature and struggle with complex literary forms. Comparing tools for Danish literary sentiment analysis, [13] found that dictionaries were outperformed by multilingual transformer models – likely because they leverage contextual attention. While multilingual transformers, such as mBERT and XLM-R [11], show promise for cross-lingual sentiment analysis in literature [13], cultural and linguistic biases inherited from English pretraining remain a concern [10, 37].
@@
    Cross-domain performance often drops, especially when applying tools trained on social media to literature, where sentiment is expressed in a distinct manner [3, 14, 33]. Literary language tends to be more omissive and implicit, relying less on emotionally charged vocabulary and more on concrete descriptions of objects and situations to evoke affect – a domain-specific mode of sentiment expression that models often fail to capture [14]. This domain-specificity also varies across genres: when using a model fine-tuned on Twitter posts, poetry shows the weakest correlation with human ratings, prose falls in the middle, and Facebook posts the strongest [14].
@@
    Historical data, marked by diachronic language change, often reduces model performance. While fine-tuned multilingual transformers show promise for older texts [1, 29, 13], challenges remain. Lexical drift – including semantic shift, word loss (e.g., thou, peradventure), changing frequencies, and temporal polarity shifts – limits sentiment inference, especially when models rely on priors from modern corpora. [^1] Diachronic sentiment analysis is particularly challenging for traditional machine learning approaches as words’ meaning and polarity change continuously, while most models require steady ground truths for training, creating somewhat artificial “museums” of words’ sentiment scores in a given historical period. In the case of temporal polarity shifts, even short-term changes can lower model performance, as learned polarities no longer reflect current usage [23].
@@
Moreover, a major drawback of recent transformer-based approaches is that, while they outperform dictionary-based tools on historical and literary data [13], they are typically optimized for trinary classification (positive, neutral, negative), limiting their usefulness for fine-grained sentiment analysis. Although model confidence scores can be repurposed for continuous output – with medium to strong correlation to human ratings [13] – the resulting distributions still cluster around the original three categories, producing what is effectively a pseudo-trinary distribution. This poses a problem for literary analysis tasks, not least sentiment arc modelling, where detrending methods to smoothen out the signal necessitate continuous scores. When sentiment scores behave in extreme ways – as they will with pseudo-trinary distributions – smoothing will tend to collapse variation toward the neutral midpoint, removing meaningful information.
@@
In this paper, we introduce a method to create continuous-scale sentiment scores that are more closely aligned with the distribution of human scores, while also mitigating language-, domain-, and historical data issues by training a model on the actual language and domain of the use case.
@@
We test this approach on English and Danish literary texts, comparing it to existing transformer-based models and popular dictionary-based tools, across both fiction and nonfiction genres. The benchmark includes both historical literary genres (e.g., 19th-century hymns) and contemporary texts (e.g., blogs), enabling us to evaluate model performance in settings that better reflect the needs of researchers working with multilingual or diachronic literary corpora. We pursue three aims: (1) to assess model performance on contemporary literary and non-literary texts; (2) to compare performance across literary genres; and (3) to evaluate models on historical and multilingual literary data. We begin by testing our approach on Fiction4 — a recent annotated fiction corpus that spans four literary genres, two languages (English and Danish), and a period from 1798 to 1965. We then validate our approach further on EmoBank, a standard sentiment analysis dataset that includes contemporary genres and a small set of fiction, to gauge the generalizability of our approach and to control for overfitting to literary data.
@@ 2.1 Data
is a text corpus manually annotated for sentiment according to the psychological Valence-Arousal-Dominance scheme. It was compiled at JULIE Lab, Jena University [9], [^2] https://github.com/JULIELab/EmoBank/, containing sentences from the MASC dataset, which is diverse both in terms of overall composition with diverse domains, and topically within categories. [^3] On some EmoBank categories: Essays includes eight texts, i.a., “A Brief History of Steel in Northeastern Ohio’. Fiction comprises six prose pieces across genres, i.a., Richard Harding’s “A Wasted Day” and the SciFi story “Captured Moments”. Newspapers contain reports (e.g., “A.L. Williams Corp. was merged into Primerica Corp.”) and longer reportages. Travel Guides are written in prose, including both place histories (e.g., “A Brief History of Jerusalem”) and reflective pieces (e.g., “Dublin and the Dubliners”). See the full MASC corpus at: https://anc.org/data/masc/corpus/browse-masc-data/. It includes six categories: Letters, Blog, Newspaper, Essays, Fiction, and Travel guides. [^4] We excluded the ‘Sem-Eval’ category as it was internally diverse. Inter Rater Reliability for the whole dataset is: Krippendorff’s $\alpha=0.34$. [^5] Since EmoBank lacks unique annotator IDs, we cannot correlate individual annotators’ scores. Instead, Krippendorff’s $\alpha$ measures agreement across ratings per item. IRR per subset is shown in Table 4. We use the mean sentence-based valence scores overall and per category to compare model performance.
@@
is a dataset of literary texts, spanning literary texts across four genres and two languages (English and Danish) in the 19th and 20th century.666 https://huggingface.co/datasets/chcaa/fiction4sentiment, for details, see [14], compiled at the Center for Humanities Computing, Aarhus University. The corpus consists of three main authors, Sylvia Plath for poetry, Ernest Hemingway for prose, and H.C. Andersen for fairytales. Hymns were collected from Danish official church hymnbooks published between 1798 and 1873. All sentences in the corpus were annotated for by at least two annotators [14]. Inter Rater Reliability for the whole dataset is: Spearman’s $\rho=0.63$ and Krippendorff’s $\alpha=0.67$. [^7] Humans rarely reach an agreement higher than 80% ($\alpha$¿0.80) for categorical tagging (positive/neutral/negative) on nonliterary texts [36] – and have lower IRR for continuous scale annotation [2] – especially of literary texts [27]. We use the mean sentence-based valence score overall, per language set, and per genre to compare model performance.
@@ 2.2.1 Dictionary-based
Because of their popularity and wide usage in literary studies [1, 5, 4], as a baseline, we tested the dictionary-based tools VADER [15] and Syuzhet [17]. They assign sentiment scores (from negative to positive) by word-score matching and specific rules. Syuzhet was developed explicitly for literary texts. [^8] The Syuzhet lexicon was developed in the Nebraska Literary Lab under the direction of Matthew L. Jockers. When using these tools, we translated Danish sentences into English. [^9] Using googletrans: https://pypi.org/project/googletrans/ A human did not review translations.
@@ 2.2.2 Transformer-based
To test transformer-based methods, we chose two multilingual models. When testing models on Danish texts, we added three models fine-tuned for Danish. These were all tested across EmoBank categories, as well as Fiction4 genres and languages. We list all models in Appendix A, Table 5.101010Code for comparing (HuggingFace-stored) sentiment models (with transformed outputs) on the Fiction4 or EmoBank is at: https://anonymous.4open.science/r/literary_sentiment_benchmarking-D6E6. One of the multilingual models – twitter-xlm – showed the best performance on Fiction4 in [13]. Danish models were picked based on their performance in a recent benchmark [20], and – in the case of MeMo-BERT-Sa – for being developed for 19th-century novels [19].
Conversion of model output: We convert Transformers’ standard three-ways outputs (positive, neutral, negative) to continuous values using their confidence scores  [^11] The score output by finetuned models (e.g., “positive”, 0.66) is a softmax-normalized class probability – a pseudo confidence score – reflecting how strongly a model prefers one label over another. It comes from the linear classification head atop models. as a proxy for intensity (e.g., positive, 0.67 → +0.67; negative → –0.67; neutral → 0). Mapping a model’s confidence values onto a continuous scale often outperforms dictionary-based tools for literary sentiment [4, 13].
@@ 2.3 Our approach
It has been claimed that concepts – such as a sentiment – are approximately represented in a linear fashion within embedding space, which is denoted by the linear representation hypothesis [24]. The hypothesis states that concepts are encoded as a direction in the embedding space and that the further you move in a given direction, the stronger the concept is represented (see Figure 1). These linear representations of semantic information have been found in both encoding and decoding models, at varying levels of abstraction [35, 34, 22, 39]. Suppose we have access to the direction that encodes sentiment. In that case, we can project any embedded sentence onto the concept vector and gauge the sentiment of any given sentence, as seen in Figure 1.
@@ 2.3.1 Concept Vector Projection
We propose an algorithm that constructs a concept vector in a given embedding space using positive and negative exemplary sentences that represent the opposing extremes of the concept. Using a pre-trained sentence embedding model $\mathbf{M}$, the algorithm embeds a labeled set of sentences $\mathbf{S}$. It assumes that a concept – here sentiment – is represented linearly in the embedding space. To define the concept vector, the algorithm computes the mean embedding of both the positive and negative sentiment examples, then calculates the vector pointing from the negative to the positive mean. This relies on the assumption that when averaging multiple sentences, all non-sentiment information will disappear as Gaussian noise with a mean of zero, leaving behind only the information encoding sentiment [18, 39].
The resulting vector then theoretically encodes sentiment direction. New sentences can be assessed for their relation to the sentiment by projecting their embeddings onto this vector: the farther along the direction the projection lies, the stronger their positive relation is. Defining the concept vector as a unit vector, the projection of a given embedding $\mathbf{e}_{i}$ onto the unit concept vector $\hat{\mathbf{v}}$ is given by the dot product: $\mathbf{e}_{i}\cdot\hat{\mathbf{v}}$. This projects the sentence embedding to the subspace spanned by the Concept Vector. The high-dimensional embedding has thereby been reduced to a one-dimensional sentiment score, as seen in figure Figure 2. Defining a concept vector requires only a set of positive and negative example sentences. This is sufficient to predict the sentiment of any subsequent sentence, whether labeled or unlabeled. The Concept Vector Projection (CVP) algorithm formally described in Appendix B. The implementation of this method is available at https://github.com/centre-for-humanities-computing/embedding-projection.

### v2

@@ Abstract
Sentiment Analysis is widely used to quantify sentiment in text, but its application to literary texts poses unique challenges due to figurative language, stylistic ambiguity, as well as sentiment evocation strategies. Traditional dictionary-based tools tend to underperform, especially for low-resource languages, and transformer models, while promising, output coarse categorical labels that limit fine-grained analysis. We introduce a novel continuous sentiment scoring method based on concept vector projection, trained on multilingual literary data, which captures nuanced sentiment expressions across genres, languages, and historical periods. Our approach outperforms existing tools on English and Danish texts, producing sentiment scores which distribution matches human ratings, improving sentiment arc modeling and analysis in literature.
@@ 1 Introduction & Related Works
Sentiment analysis quantifies sentiment in text and is widespread across domains, from product reviews analysis to social media monitoring [tsao_asymmetric_2018, bollen_twitter_2011]. Computational literary studies have employed sentiment analysis to model narrative dynamics, particularly sentiment arcs, across novels [jockers_novel_2014, reagan_emotional_2016, cellier_prediction_2016, bizzoni_sentimental_2023]. This requires continuous sentiment scores, mapping sentiment along a spectrum rather than using categorical labels like positive/negative.
Despite the growing use of continuous sentiment scoring in literary studies, the validity of current tools in capturing literary sentiment expression remains underexplored. Popular tools such as Syuzhet have faced severe criticism for oversimplification or poor generalizability [swafford2015problems] – issues that point to broader limitations in applying off-the-shelf sentiment tools to literary texts. Indeed, the literary domain poses distinct challenges: figurative language, multiple narrative layers, and stylistic ambiguity all complicate sentiment detection [booth_rhetoric_1983, bizzoni_feldkamp_sea_2024].
More recent transformer-based models appear better equipped to handle the complexity of literary language [schmidt_using_2021], and techniques exist to transform categorical model outputs into continuous scores [bizzoni_comparing_2023]. This method has proven more effective than tailored dictionary-based tools, particularly in low-resource language settings and across languages [feldkamp_comparing_2024]. However, empirical benchmarks comparing model predictions to human judgments remain limited in languages other than English.
@@
    Cross-lingual performance drops. Most Sentiment Analysis tools tackle high-resource languages, and their transfer to low-resource ones like Danish is non-trivial. Although Danish has several dictionary-based tools (i.a., Asent [Enevoldsen_Asent_Fast_flexible_2022], Sentida [lauridsen_sentida_2019]), these have seen little use on historical literature and struggle with complex literary forms. Comparing tools for Danish literary sentiment analysis, [feldkamp_comparing_2024] found that multilingual transformer models outperfromed dictionaries – likely because they leverage contextual attention. While multilingual transformers, such as mBERT and XLM-R [Devlin_2019], show promise for cross-lingual sentiment analysis in literature [feldkamp_comparing_2024], cultural and linguistic biases inherited from English pretraining remain a concern [de_bruyne_2022_language, xu_survey_2022].
@@
    Cross-domain performance often drops, especially when applying tools trained on social media to literature, where sentiment is expressed in a distinct and complex manner [bizzoni_feldkamp_sea_2024, feldkamp2024sentiment, vishnubhotla-etal-2024-emotion]. Literary language tends to be more omissive and implicit, relying less on charged vocabulary and more on concrete descriptions of objects and situations to evoke affect – a domain-specific mode of sentiment expression that models fail to capture [feldkamp2024sentiment]. This domain-specificity varies across domains: when using a model fine-tuned on Twitter posts, poetry shows the weakest correlation with human ratings, prose falls in the middle, and Facebook posts show the strongest correlation [feldkamp2024sentiment].
@@
    Historical data, marked by diachronic language change, reduces model performance. While fine-tuned multilingual transformers show promise [allaith_sentiment_2023, schmidt_evaluation_2018, feldkamp_comparing_2024], challenges remain. Lexical drift – including semantic shift, word loss (e.g., thou, peradventure), changing frequencies, and temporal polarity shifts – limits sentiment inference if models rely on priors from modern corpora. [^1] Diachronic sentiment analysis is challenging for traditional machine learning approaches as words’ meaning and polarity change in a continuous way, while most models require steady ground truths for training, creating artificial “museums” of words’ sentiment scores in a given historical period. For temporal polarity shifts, even short-term changes can lower model performance [lukes_sentiment_2018].
@@
A major drawback of recent transformer-based approaches is that, while they outperform dictionary-based tools on historical and literary data [feldkamp_comparing_2024], they tend to perform trinary classifications (positive, neutral, negative), limiting their usefulness for fine-grained sentiment analysis. Although model confidence scores can be repurposed for continuous output – with medium to strong correlation to human ratings [feldkamp_comparing_2024] – the resulting distributions still cluster around the original three categories, producing what is effectively a pseudo-trinary distribution. This poses a problem for literary analysis tasks, not least sentiment arc modelling, where detrending methods to smoothen out the signal necessitate continuous scores. When sentiment scores behave in extreme ways – as they will with pseudo-trinary distributions – smoothing will tend to collapse variation toward the neutral midpoint, removing meaningful information.
@@
In this paper, we introduce a method to create continuous-scale sentiment scores that are more closely aligned with the distribution of human scores, while also mitigating language-, domain-, and historical data issues by basing the method on the language and domain of the use case.
@@
We test this approach on English and Danish literary texts, comparing it to existing transformer-based models and popular dictionary-based tools, across both fiction and nonfiction genres. The benchmark includes both historical literary genres (e.g., 19th-century hymns) and contemporary texts (e.g., blogs), enabling us to evaluate model performance in settings that better reflect the needs of researchers working with multilingual or diachronic literary corpora. We pursue three aims: (1) to assess model performance on contemporary literary and non-literary texts; (2) to compare performance across literary genres; and (3) to evaluate models on historical and multilingual literary data. We begin by testing our approach on Fiction4 — a recent annotated fiction corpus that spans four literary genres, two languages (English and Danish), in the period 1798 to 1965. We then validate our approach further on EmoBank, a standard sentiment analysis dataset that includes contemporary genres and a small set of fiction, to gauge the generalizability of our approach and to control for overfitting to literary data.
@@ 2.1 Data
is a text corpus manually annotated for sentiment according to the psychological Valence-Arousal-Dominance scheme. It was compiled at JULIE Lab, Jena University [buechel-hahn-2017-emobank], [^2] https://github.com/JULIELab/EmoBank/, containing sentences from the MASC dataset, which is diverse both in terms of overall composition with diverse domains, and topically within categories. [^3] On some EmoBank categories: Essays includes eight texts, i.a., “A Brief History of Steel in Northeastern Ohio’. Fiction comprises six prose pieces across genres, i.a., Richard Harding’s “A Wasted Day” and the SciFi story “Captured Moments”. Newspapers contain reports (e.g., “A.L. Williams Corp. was merged into Primerica Corp.”) and longer reportages. Travel Guides are written in prose, including both place histories (e.g., “A Brief History of Jerusalem”) and reflective pieces (e.g., “Dublin and the Dubliners”). See the full MASC corpus at: https://anc.org/data/masc/corpus/browse-masc-data/. It includes six categories: Letters, Blog, Newspaper, Essays, Fiction, and Travel guides. [^4] We excluded the ‘Sem-Eval’ category as it was internally diverse. Inter Rater Reliability for the whole dataset is: Krippendorff’s $\alpha=0.34$. [^5] Since EmoBank lacks unique annotator IDs, we cannot correlate individual annotators’ scores. Instead, Krippendorff’s $\alpha$ measures agreement across ratings per item. IRR per subset is shown in Table 4. We use the mean sentence-based valence scores overall and per category to compare model performance.
@@
is a dataset of literary texts, spanning literary texts across four genres and two languages (English and Danish) in the 19th and 20th century. [^6] https://huggingface.co/datasets/chcaa/fiction4sentiment, for details, see [feldkamp2024sentiment], compiled at the Center for Humanities Computing, Aarhus University. The corpus consists of three main authors, Sylvia Plath for poetry, Ernest Hemingway for prose, and H.C. Andersen for fairytales. Hymns were collected from Danish official church hymnbooks published between 1798 and 1873. All sentences in the corpus were annotated for by at least two annotators [feldkamp2024sentiment]. Inter Rater Reliability for the whole dataset is: Spearman’s $\rho=0.63$ and Krippendorff’s $\alpha=0.67$. [^7] Humans rarely reach an agreement higher than 80% ($\alpha$¿0.80) for categorical tagging (positive/neutral/negative) on nonliterary texts [wilson_recognizing_2005] – and have lower IRR for continuous scale annotation [batanovic_versatile_2020] – especially of literary texts [rebora_comparing_2023]. We use the mean sentence-based valence score overall, per language set, and per genre to compare model performance.
@@ 2.2.1 Dictionary-based
Because of their popularity and wide usage in literary studies [allaith_sentiment_2023, bizzoni_nobels, bizzoni_comparing_2023], as a baseline, we tested the dictionary-based tools VADER [hutto_vader_2014] and Syuzhet [syuzhet]. They assign sentiment scores (from negative to positive) by word-score matching and specific rules. Syuzhet was developed explicitly for literary texts. [^8] The Syuzhet lexicon was developed in the Nebraska Literary Lab under the direction of Matthew L. Jockers. When using these tools, we translated Danish sentences into English as they do not perform well on the original Danish. [^9] Using googletrans: https://pypi.org/project/googletrans/. Humans did not review translations. As such, dictionaries represent a rough baseline.
@@ 2.2.2 Transformer-based
To test transformer-based methods, we chose two multilingual models. When testing models on Danish texts, we added three models fine-tuned for Danish. These were all tested across EmoBank categories, as well as Fiction4 genres and languages. We list all models in Appendix A, Table 5.101010Code for comparing (HuggingFace-stored) sentiment models (with transformed outputs) on the Fiction4 or EmoBank is at: https://github.com/centre-for-humanities-computing/literary_sentiment_benchmarking. One of the multilingual models – twitter-xlm – showed the best performance on Fiction4 in [feldkamp_comparing_2024]. Danish models were picked based on their performance in a recent benchmark [allaith-etal-2023-sentiment], and – in the case of MeMo-BERT-Sa – for being developed for 19th-century novels [al-laith_development_2024].
Conversion of model output: We convert Transformers’ standard three-ways outputs (positive, neutral, negative) to continuous values using their confidence scores  [^11] The score output by finetuned models (e.g., “positive”, 0.66) is a softmax-normalized class probability – a pseudo confidence score – reflecting how strongly a model prefers one label over another. It comes from the linear classification head atop models. as a proxy for intensity (e.g., positive, 0.67 → +0.67; negative → –0.67; neutral → 0). Mapping a model’s confidence values to a continuous scale often outperforms dictionary-based tools for literary sentiment [bizzoni_comparing_2023, feldkamp_comparing_2024].
@@
We tested transformer-based models on the original Danish and English, as well as on the Danish sentences translated to English (see Table 3), since one study found that some models work better on google-translated sentences [feldkamp_comparing_2024], perhaps as the translation acts to standardize historic forms.
@@ 2.3 Our approach
It has been claimed that concepts – such as a sentiment – are approximately represented in a linear fashion within embedding space, which is denoted by the linear representation hypothesis [park_linear_2024]. The hypothesis states that concepts are encoded as a direction in the embedding space and that the further you move in a given direction, the stronger the concept is represented (see Figure 1). These linear representations of semantic information have been found in both encoding and decoding models, at varying levels of abstraction [wehner_taxonomy_2025, vu-parker-2016-k, li-etal-2021-implicit, zhao2024beyond]. Suppose we have access to the direction that encodes sentiment. In that case, we can project any embedded sentence onto the concept vector and gauge the sentiment of any given sentence, as seen in Figure 1.
@@ 2.3.1 Concept Vector Projection
We propose an algorithm that constructs a concept vector in a given embedding space using positive and negative exemplary sentences that represent the opposing extremes of the concept. Using a pre-trained sentence embedding model $\mathbf{M}$, the algorithm embeds a labeled set of sentences $\mathbf{S}$. It assumes that a concept – here sentiment – is represented linearly in the embedding space. To define the concept vector, the algorithm computes the mean embedding of both the positive and negative sentiment examples, then calculates the vector pointing from the negative to the positive mean. This relies on the assumption that when averaging multiple sentences, all non-sentiment information will disappear as Gaussian noise with a mean of zero, leaving behind only the information encoding sentiment [kim2018interpretability, zhao2024beyond].
The resulting vector then theoretically encodes sentiment direction. New sentences can be assessed for their relation to the sentiment by projecting their embeddings onto this vector: the farther along the direction the projection lies, the stronger their positive relation is. Defining the concept vector as a unit vector, the projection of a given embedding $\mathbf{e}_{i}$ onto the unit concept vector $\hat{\mathbf{v}}$ is given by the dot product: $\mathbf{e}_{i}\cdot\hat{\mathbf{v}}$. This projects the sentence embedding to the subspace spanned by the Concept Vector. The high-dimensional embedding has thereby been reduced to a one-dimensional sentiment score, as seen in figure Figure 2. Defining a concept vector requires only a set of positive and negative example sentences. This suffices to predict the sentiment of any subsequent sentence, whether labeled or unlabeled. The Concept Vector Projection (CVP) algorithm formally described in Appendix C. The implementation of this method is available at https://github.com/centre-for-humanities-computing/embedding-projection.